OpenStudy (anonymous):

Check my answer trig function values Find the remaining five trigonometric function values Sec θ =-8 and θ is in the second quadrant sin θ= radical 63/8 cos θ = -1/8 tan θ = -radical 63 cot θ = -radical 63/radical63 csc θ = 8radical63/63

OpenStudy (anonymous):

find the other side via pythagoras, then you can find all ratios

OpenStudy (anonymous):

isnt the hypotenuse 8?

OpenStudy (anonymous):



OpenStudy (anonymous):

x is -1?

OpenStudy (anonymous):

first one is right second one is right

OpenStudy (anonymous):

yeah i just labelled it as a length you are in quadrant 2 so cosine is negative and sine is positive etc

OpenStudy (anonymous):

third one is right

OpenStudy (anonymous):

cotangent is wrong

OpenStudy (anonymous):

reciprocal of tangent, should be \(-\frac{1}{\sqrt{63}}\)

OpenStudy (anonymous):

cosecant is right

OpenStudy (anonymous):

but you can't have radical in the denominator

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then it is \[-\frac{\sqrt{63}}{63}\]

OpenStudy (anonymous):

k ty

OpenStudy (anonymous):

yw